OpenStudy (anonymous):

The number of passengers of terminal A, A(n) and terminal B, B(n), in the nth year are \(A(n)=2pq^n\) and \(B(n) = npq^2\), where p and q are positive constants. Show that the number of passenger of terminal B will not exceed that of terminal A in the future.

OpenStudy (anonymous):

Sorry but can you explain constants i never get the parts of a equation right and no more than i know positive constants

OpenStudy (anonymous):

Im good at equations just not the parts at all

OpenStudy (anonymous):

Hmm, let's say, it is given that p=2000, q=495.

OpenStudy (anonymous):

Sorry but im still dumbfounded let me get my friend

OpenStudy (anonymous):

@kewlgeek555

OpenStudy (anonymous):

Give me some probability no matter how hard i can do it in time but sometimes some equations just stupify me

OpenStudy (kewlgeek555):

(o-o) I don't think I am this advanced in Algebra, yet. (^_^;)

OpenStudy (anonymous):

If number of passenger of terminal B will exceed that of terminal A, then \[npq^2>2pq^n\] for some positive integers n.

OpenStudy (anonymous):

Yeah I can do this but it would probably just be quicker to get your local tutor service or a teacher to help. Try getting some of the higher level Helpers to get this by typing: @TheirName

OpenStudy (anonymous):

LOL, I was tutoring when I overcame this question!

OpenStudy (anonymous):

I can do this i learned it this year but the way i know 1 Number and 1 Exponent on different sides isnt enough for me to solve the... I know what it is called but i cant remember its a Equation but with larger smaller equal to equal to or less equal to or larger

OpenStudy (anonymous):

Could you give me some numbers for some of the variables so that it would still work and still be solvable

OpenStudy (anonymous):

and it would be \[npq^{2}\ge2pq ^{n}\]

OpenStudy (anonymous):

Does not exceed it does not say cannot be equal to

OpenStudy (anonymous):

"the number of passenger of terminal B will not exceed that of terminal A " B(n) <= A(n) :)

OpenStudy (anonymous):

So if you could solve atleast one var on each side i can solve it

OpenStudy (anonymous):

You mixed up the two function, perhaps :)

OpenStudy (anonymous):

SO could you solve some vars please or make some that work

OpenStudy (anonymous):

If I could solve it, I would not have posted it here :)

OpenStudy (anonymous):

Ok then let me think

OpenStudy (anonymous):

Im guessing that Npq stands for B(here) amd 2pq stands for A(here)

OpenStudy (anonymous):

Well, \(B(n)=npq^2\) is a function in n.

OpenStudy (anonymous):

JUST GO USE A INTERNET ALGEBRA CALCULATOR IM SORRY BUT I JUST CANT

OpenStudy (anonymous):

Such calculators are not allowed in the exams...

OpenStudy (anonymous):

(-_-) |____| \ \ [=] / / / (@_@)// / / / I RAGE QUIT THIS MATH!!!!

OpenStudy (anonymous):

(-_-) /| |\ |____| [=] // (-_-) // / | \|\ / / / I JUST QUIT

OpenStudy (anonymous):

I CANT DO IT

OpenStudy (anonymous):

I TRIED AND FAILED BUT ATLEAST I TRIED

OpenStudy (anonymous):

In one line A grows exponentially with n and B grows linearly with n. After a while, A will always be bigger than B \[ A(n)=C_1 q^n\\ B(n)=C_2 n \]

OpenStudy (anonymous):

We have to suppose that q>1,

OpenStudy (anonymous):

That's what I have told the girl I tutored. However, in case we must use a more "mathematical" way to do it, what is the possible approach to solve this problem?

OpenStudy (anonymous):

Show her that if q>1, then \[ \lim_{n\to \infty}\frac{A(n)}{B(n)}=\infty \] Hence there is N, so n>N \[ \frac{A(n)}{B(n)}>1\\ A(n) > B(n) \]

OpenStudy (anonymous):

I don't mind explaining the concept of limit to her, since she has not learnt it. Though, I bet she is not interested since it is out of syllabus. Thanks for your suggestions. :)